THE ABUNDANCE OF C_{7}^{−} IN DIFFUSE INTERSTELLAR CLOUDS

\underline{D. P. RUFFLE}, R. TERZIEVA, E. HERBST, Departments of Physics and Astronomy, and Chemical Physics Program, The Ohio State University, Columbus, OH 43210.

It has recently been proposed that C_{7}^{−} is a carrier of five of the numerous diffuse interstellar bands. In this talk, we report an attempt to reproduce the required abundance of C_{7}^{−} relative to hydrogen of 4 × 10^{−8} through the use of various models for the chemistry in diffuse clouds. Our models incorporate large neutral and charged linear hydrocarbons, and contain the time-dependent desorption of seed molecules from grain surfaces. We are able to find conditions that lead to the production of the required fractional abundance of C_{7}^{−} but only for short periods of time in an astronomical sense.